# Hyper-SD: Trajectory Segmented Consistency Model for Efficient Image Synthesis

### v1

@@ Abstract
Figure 1: The visual comparison between our Hyper-SDXL and other methods. From the first to the fourth column, the prompts for these images are (1) a dog wearing a white t-shirt, with the word “hyper” written on it … (2) abstract beauty, approaching perfection, pure form, golden ratio, minimalistic, unfinished,… (3) a crystal heart laying on moss in a serene zen garden … (4) anthropomorphic art of a scientist stag, victorian inspired clothing by krenz cushart …., respectively.
11footnotetext: Project Lead.00footnotetext: Correspondence to: xiaoxuefeng.ailab@bytedance.com
@@ 1 Introduction
To navigate these hurdles and harness the full potential of DMs, we present an advanced framework that adeptly combines trajectory-preserving and trajectory-reformulating distillation techniques. Firstly, we proposed trajectory segmented consistency distillation (TSCD), which divides the time steps into segments and enforces consistency within each segment while gradually reducing the number of segments to achieve all-time consistency. This approach addresses the issue of suboptimal consistency model performance caused by insufficient model fitting capability and accumulated errors in inference. Secondly, we leverage human feedback learning techniques [28, 32, 17] to optimize the accelerated model, modifying the ODE trajectories to better suit few-step inference. This results in significant performance improvements, even surpassing the capabilities of the original model in some scenarios. Thirdly, we enhanced the one-step generation performance using score distillation [26, 31], achieving the idealized all-time consistent model via a unified LORA. In summary, our main contributions are summarized as follows:
  • Accelerate: we propose TSCD that achieves a more fine-grained and high-order consistency distillation approach for the original score-based model.
  • Boost: we incorpoate human feedback learning to further enhance model performance in low-steps regime.
  • Unify: we provide a unified LORA as the all-time consistency model and support inference at all NTEs.
  • Performance: Hyper-SD achieves SOTA performance in low-steps inference for both SDXL and SD1.5.
@@ 2 Preliminaries
2.1 Diffusion Model
$\mathrm{d}x_{t}=\mu(x_{t},t)\mathrm{d}t+\sigma(t)\mathrm{d}w_{t},$ (1)
@@ 3.1 Trajectory Segmented Consistency Distillation
$L_{TSCD}=d($ $\Psi(x_{t_{n}},f_{stu}(x_{t_{n}},t_{n},c),t_{end}),$ (14)
$\Psi(\hat{x}_{t_{n-1}},f^{-}_{stu}(\hat{x}_{t_{n-1}},t_{n-1},c),t_{end}))$
where $\hat{x}_{t_{n-1}}$ is computed as Eq. 3, and $f^{-}_{stu}$ denotes the Exponential Moving Average (EMA) of $f_{stu}$.
@@
The TSCD method offers two principal advantages: Firstly, fine-grained segment distillation reduces model fitting complexity and minimizes errors, thus mitigating degradation in generation quality. Secondly, it ensures the preservation of the original ODE trajectory. Models from each training stage can be utilized for inference at corresponding steps while closely mirroring the original model’s generation quality. We illustrate the complete procedure of Progressive Consistency Distillation in Algorithm 1. It is worth noting that, by utilizing Low-Rank Adaptation(LoRA) technology, we train TSCD models as plug-and-play plugins that can be used instantly.
@@
3.2 Human Feedback Learning
3.3 One-step Generation Enhancement
4 Experiments
4.1 Implementation Details
Dataset. We use a subset of the LAION and COYO datasets following SDXL-lightning [8] during the training procedure of Sec 3.1 and Sec 3.3. For the Human Feedback Learning in Sec 3.2, we generated approximately 140k artist-style text images for style optimization using the SDXL-Base model and utilized the COCO2017 train split dataset with instance annotations and captions for structure optimization.
Training Setting. For TSCD in Sec 3.1, we progressively reduced the time-steps segments number as $8\rightarrow 4\rightarrow 2\rightarrow 1$ in four stages, employing 512 batch size and learning rate $61e-$ 6 across 32 NVIDIA A100 80GB GPUs. We trained Lora instead of Unet for all the distillation stages for convenience, and the corresponding Lora is loaded to process the human feedback learning optimization in Sec 3.2. For one-step enhancement in Sec 3.3, we trained the unified all-timesteps consistency Lora with time-step inputs $T=999$ and the dedicated model for single-step generation with $T=800$.
Evaluation Metrics. We use the aesthetic predictor pre-trained on the LAION dataset and CLIP score(ViT-B/32) to evaluate the visual appeal of the generated image and the text-to-image alignment. We further include some recently proposed metrics, such as ImageReward score [28], and Pickscore [5] to offer a more comprehensive evaluation of the model performance. Note that we do not report the Fréchet Inception Distance(FID) as we observe it can not well reflect the actual generated image quality in our experiments. In addition to these, due to the inherently subjective nature of image generation evaluation, we conduct an extensive user study to evaluate the performance more accurately.
Method Steps Support Arch. CFG Free One-Step UNet One-Step LoRA
PeRFlow [29] 4+ SD15 No No No
TCD [34] 2+ SD15/XL Yes No No
LCM [12] 2+ SD15/XL Yes No No
Turbo [21] 1+ SD21/XL Yes Yes No
Lightning [8] 1+ SDXL Yes Yes No
Ours 1+ SD15/XL Yes Yes Yes
Table 1: Comparison with other acceleration approaches.
Model Steps Type CLIP Score Aes Score Image Reward Pick Score
SD15-Base [18] 25 UNet 31.88 5.26 0.18 0.217
SD15-LCM [12] 4 LoRA 30.36 5.66 -0.37 0.212
SD15-TCD [34] 4 LoRA 30.62 5.45 -0.15 0.214
PeRFlow [29] 4 UNet 30.77 5.64 -0.35 0.208
Hyper-SD15 1 LoRA 30.87 5.79 0.29 0.215
SDXL-Base [18] 25 UNet 33.16 5.54 0.87 0.229
SDXL-LCM [12] 4 LoRA 32.43 5.42 0.48 0.224
SDXL-TCD [34] 4 LoRA 32.45 5.42 0.67 0.226
SDXL-Lightning [8] 4 LoRA 32.40 5.63 0.72 0.229
Hyper-SDXL 4 LoRA 32.56 5.74 0.93 0.232
SDXL-Turbo [21] 1 UNet 32.33 5.33 0.78 0.228
SDXL-Lightning [8] 1 UNet 32.17 5.34 0.54 0.223
Hyper-SDXL 1 UNet 32.85 5.85 1.19 0.231
Refer to captionRefer to caption
@@ 4.1 Implementation Details
4.2 Main Results
Qualitative Comparison. In Figs. 5, 4 and 6, we present comprehensive visual comparison with recent approaches, including LCM [12], TCD [34], PeRFLow [29], Turbo [21] and Lightning [8]. Our observations can be summarized as follows. (1) Thanks to the fact that SDXL has almost 2.6B parameters, the model is able to synthesis decent images in 4 steps after different distillation algorithms. Our method further utilizes its huge model capacity to compress the number of steps required for high-quality outcomes to 1 step only, and far outperforms other methods in terms of style (a), aesthetics (b-c) and image-text matching (d) as indicated in Fig. 4. (2) On the contrary, limited by the capacity of SD15 model, the images generated by other approaches tend to exhibit severe quality degradation. While our Hyper-SD consistently yields better results across different types of user prompts, including photographic (a), realistic (b-c) and artstyles (d) as depicted in Fig. 5. (3) To further release the potential of our methodology, we also conduct experiments on the fully fine-tuning of SDXL model following previous works [8, 21]. As shown in Fig. 6, our 1-Step UNet again demonstrates superior generation quality that far exceeds the rest of the opponents. Both in terms of colorization (a-b) and details (c-d), our images are more presentable and attractive when it comes to the real-world application scenarios.
@@ 4.2 Main Results
Arch. Steps CLIP Score Aes Score Image Reward Pick Score
SD15 512px 8 30.73 5.47 0.53 0.224
4 31.07 5.55 0.53 0.224
2 31.21 5.93 0.45 0.222
1 30.87 5.79 0.29 0.215
SDXL 1024px 8 32.54 5.83 1.14 0.233
4 32.51 5.52 1.15 0.234
2 32.59 5.71 1.15 0.234
1 32.59 5.69 1.06 0.226
Table 3: Quantitative results on unified LoRAs.
4.3 Ablation Study
Unified LoRA. In addition to the different steps of LoRAs proposed above, we note that our one-step LoRA can be considered as a unified approach, since it is able to reason about different number of steps (e.g. 1,2,4,8 as shown in Fig. 8) and consistently generate high-quality results under the effect of consistency distillation. For completeness, Tab. 3 also presents the quantitative results of different steps when applying the 1-Step unified LoRA. We can observe that there is no difference in image-text matching between different steps as the CLIPScore evaluates, which means that user prompts are well adhered to. And as the other metrics show, the aesthetics rise slightly as the step increases, which is as expected after all the user can choose based on the needs for efficiency. This would be of great convenience and practicality in real-world deployment scenarios, since generally only one model can be loaded per instance.
Compatibility with Base Model. Fig. 9 shows that our LoRAs can be applied to different base models. Specifically, we conduct experiments on anime [^1] https://civitai.com/models/112902, realistic [^2] https://civitai.com/models/133005 and artstyle [^3] https://civitai.com/models/119229 base models. The results demonstrate that our method has a wide range of applications, and the lightweight LoRA also significantly reduces the cost of acceleration.
Compatibility with ControlNet. Fig. 10 shows that our models are also compatible with ControlNet [33]. We test the one-step unified SD15 and SDXL LoRAs on the scribble [^4] https://huggingface.co/lllyasviel/control_v11p_sd15_scribble and canny [^5] https://huggingface.co/diffusers/controlnet-canny-sdxl-1.0 control images, respectively. And we can observe the conditions are well followed and the consistency of our unified LoRAs can still be demonstrated, where the quality of generated images under different inference steps are always guaranteed.
5 Discussion and Limitation
@@
6 Conclusion
We propose Hyper-SD, a unified framework that maximizes the few-step generation capacity of diffusion models, achieving new SOTA performance based on SDXL and SD15. By employing trajectory-segmented consistency distillation, we enhanced the trajectory preservation ability during distillation, approaching the generation proficiency of the original model. Then, human feedback learning and variational score distillation stimulated the potential for few-step inference, resulting in a more optimal and efficient trajectory for generating models. We have open-sourced Lora plugins for SDXL and SD15 from 1 to 8 steps inference, along with a dedicated one-step SDXL model, aiming to further propel the development of the generative AI community.

### v2

@@ Abstract
Figure 1: The visual comparison between our Hyper-SDXL and other methods. From the first to the fourth column, the prompts for these images are (1) a dog wearing a white t-shirt, with the word “hyper" written on it … (2) abstract beauty, approaching perfection, pure form, golden ratio, minimalistic, unfinished,… (3) a crystal heart laying on moss in a serene zen garden … (4) anthropomorphic art of a scientist stag, victorian inspired clothing by krenz cushart …., respectively.
@@ 1 Introduction
To navigate these hurdles and harness the full potential of DMs, we present an advanced framework that adeptly combines trajectory-preserving and trajectory-reformulating distillation techniques. Firstly, we proposed trajectory segmented consistency distillation (TSCD), which divides the time steps into segments and enforces consistency within each segment while gradually reducing the number of segments to achieve all-time consistency. This approach addresses the issue of suboptimal consistency model performance caused by insufficient model fitting capability and accumulated errors in inference. Secondly, we leverage human feedback learning techniques [35, 39, 22] to optimize the accelerated model, modifying the ODE trajectories to better suit few-step inference. This results in significant performance improvements, even surpassing the capabilities of the original model in some scenarios. Thirdly, we enhanced the one-step generation performance using score distillation [33, 38], achieving the idealized all-time consistent model via a unified LoRA.
In summary, our main contributions are summarized as follows: 1) Accelerate: we propose TSCD that achieves a more fine-grained and high-order consistency distillation approach for the original score-based model. 2) Boost: we incorpoate human feedback learning to further enhance model performance in low-steps regime. 3) Unify: we provide a unified LORA as the all-time consistency model and support inference at all NTEs. 4) Performance: Hyper-SD achieves SOTA performance in low-steps inference for both SDXL and SD1.5.
@@ 2 Preliminaries
For completeness, the preliminaries on diffusion model are provided in Appendix A.
@@ 3.1 Trajectory Segmented Consistency Distillation
$L_{TSCD}=d(\Psi(x_{t_{n}},f_{stu}(x_{t_{n}},t_{n},c),t_{end}),\Psi(\hat{x}_{t_{n-1}},f^{-}_{stu}(\hat{x}_{t_{n-1}},t_{n-1},c),t_{end}))$ (12)
where $\hat{x}_{t_{n-1}}$ refers to Eq. 1, and $f^{-}_{stu}$ denotes the Exponential Moving Average (EMA) of $f_{stu}$.
@@
The TSCD method offers two principal advantages: Firstly, fine-grained segment distillation reduces model fitting complexity and minimizes errors, thus mitigating degradation in generation quality. Secondly, it ensures the preservation of the original ODE trajectory. Models from each training stage can be utilized for inference at corresponding steps while closely mirroring the original model’s generation quality. We illustrate the complete procedure of Trajectory Segmented Consistency Distillation in Appendix B. It is worth noting that, by utilizing Low-Rank Adaptation(LoRA) technology, we train TSCD models as plug-and-play plugins that can be used instantly.
3.2 Human Feedback Learning
3.3 One-step Generation Enhancement
4 Experiments
4.1 Implementation Details
Table 1: Comparison with other acceleration approaches.
Method Step Support Arch. CFG Free One-Step UNet One-Step LoRA
PeRFlow [36] 4+ SD15 No No No
TCD [41] 2+ SD15/XL Yes No No
LCM [18] 2+ SD15/XL Yes No No
Turbo [28] 1+ SD21/XL Yes Yes No
Lightning [14] 1+ SDXL Yes Yes No
Ours 1+ SD15/XL Yes Yes Yes
Dataset. We use a subset of the LAION [29] and COYO [6] datasets following SDXL-lightning [14] during the training procedure of Sec 3.1 and Sec 3.3. For the Human Feedback Learning in Sec 3.2, we utilized the COCO2017 train split dataset with instance annotations and captions for structure optimization.
Training Setting. For TSCD in Sec 3.1, we progressively reduced the time-steps segments number as $8\rightarrow 4\rightarrow 2\rightarrow 1$ in four stages, employing 512 batch size and learning rate $61e-$ 6. We take the SOLO [32] as the instance segmentation model to achieve feedback learning in Sec 3.2. Our training per stage costs around 200 A100 GPU hours. We trained LoRA instead of UNet for all the distillation stages for convenience, and the corresponding LoRA is loaded to process the human feedback learning optimization in Sec 3.2. For one-step enhancement in Sec 3.3, we trained the unified all-timesteps consistency LoRA with time-step inputs $T=999$ and the dedicated model for single-step generation with $T=800$.
Evaluation Metrics. We use the aesthetic predictor pre-trained on the LAION dataset and CLIP score(ViT-B/32) to evaluate the visual appeal of the generated image and the text-to-image alignment. We further include some recently proposed metrics, such as ImageReward score [35], and Pickscore [11] to offer a more comprehensive evaluation of the model performance. In addition to these, due to the inherently subjective nature of image generation evaluation, we conduct an extensive user study to evaluate the performance more accurately.
4.2 Main Results
Model Step Type CLIP Score Aes Score Image Reward Pick Score
SD15-Base [24] 25 UNet 31.88 5.26 0.18 0.217
SD15-LCM [18] 4 LoRA 30.36 5.66 -0.37 0.212
SD15-TCD [41] 4 LoRA 30.62 5.45 -0.15 0.214
PeRFlow [36] 4 UNet 30.77 5.64 -0.35 0.208
Hyper-SD15 1 LoRA 30.87 5.79 0.29 0.215
SDXL-Base [24] 25 UNet 33.16 5.54 0.87 0.229
SDXL-LCM [18] 4 LoRA 32.43 5.42 0.48 0.224
SDXL-TCD [41] 4 LoRA 32.45 5.42 0.67 0.226
SDXL-Lightning [14] 4 LoRA 32.40 5.63 0.72 0.229
Hyper-SDXL 4 LoRA 32.56 5.74 0.93 0.232
SDXL-Turbo [28] 1 UNet 32.33 5.33 0.78 0.228
SDXL-Lightning [14] 1 UNet 32.17 5.34 0.54 0.223
Hyper-SDXL 1 UNet 32.85 5.85 1.19 0.231
Refer to captionRefer to caption
Qualitative Comparison. We present comprehensive visual comparison with recent approaches, including LCM [18], TCD [41], PeRFLow [36], Turbo [28] and Lightning [14]. Our observations can be summarized as follows. (1) Thanks to the fact that SDXL has almost 2.6B parameters, the model is able to synthesis decent images in 4 steps after different distillation algorithms. Our method further utilizes its huge model capacity to compress the number of steps required for high-quality outcomes to 1 step only, and far outperforms other methods in terms of style (a), aesthetics (b-c) and image-text matching (d) as indicated in Fig. 4. (2) On the contrary, limited by the capacity of SD15 model, the images generated by other approaches tend to exhibit severe quality degradation. While our Hyper-SD consistently yields better results across different types of user prompts, including photographic (a), realistic (b-c) and artstyles (d) as depicted in Sec. C.1. (3) To further release the potential of our methodology, we also conduct experiments on the fully fine-tuning of SDXL model following previous works [14, 28]. As shown in Sec. C.2, our 1-Step UNet again demonstrates superior generation quality that far exceeds the rest of the opponents. Both in terms of colorization (a-b) and details (c-d), our images are more presentable and attractive when it comes to the real-world application scenarios.
4.3 Ablation Study
In Tab. 4, we provide ablation studies on TSCD and human feedback with quantitative evaluation.
Effect of TSCD. Without the utilization of human feedback, the results show that our proposed TSCD outperforms the baseline TCD [41] significantly when inference step is lower, while the performance approaches the same as the step increases. This demonstrates the effectiveness of our progressive strategy that alleviates the training difficulties when the step is extremely low.
Effect of Human Feedback. To verify the benefit of incorporating reward models, we also conduct experiments ablating on human feedback learning of different steps. As shown in Tab. 4, the performance degradation caused by distillation process is well compensated after human feedback learning. Moreover, the image-text alignment (i.e. CLIPScore) and aesthetics (i.e. Others) evaluated on different steps are very similar, which better matches the definition of consistency model [30].
Table 3: Ablation studies of TSCD and human feedback.
Method Step CLIP Score Aes Score Image Reward Pick Score
SDXL Architecture (w/o Human Feedback)
TCD [41] 2 32.36 5.62 0.29 0.217
TSCD (Ours) 2 32.49 5.58 0.64 0.222
TCD [41] 4 32.45 5.42 0.67 0.226
TSCD (Ours) 4 32.53 5.66 0.78 0.229
TCD [41] 8 32.47 5.78 0.76 0.229
TSCD (Ours) 8 32.46 5.85 0.77 0.229
SDXL Architecture (w/ Human Feedback)
DMD [38] 1 32.70 5.58 0.82 0.223
TSCD (Ours) 1 32.59 5.69 1.06 0.226
TSCD (Ours) 2 32.61 5.84 1.04 0.232
TSCD (Ours) 4 32.56 5.74 0.93 0.232
TSCD (Ours) 8 32.56 5.89 0.93 0.232
Table 4: Quantitative results on unified LoRAs.
Arch. Step CLIP Score Aes Score Image Reward Pick Score
SD15 512px 8 30.73 5.47 0.53 0.224
4 31.07 5.55 0.53 0.224
2 31.21 5.93 0.45 0.222
1 30.87 5.79 0.29 0.215
SDXL 1024px 8 32.54 5.83 1.14 0.233
4 32.51 5.52 1.15 0.234
2 32.59 5.71 1.15 0.234
1 32.59 5.69 1.06 0.226
Effect of One-Step Enhancement. In Tab. 4, we also re-implement the DMD [38] with human feedback. The results show that our TSCD model indeed exhibit less estimation error than the original score-based model with similar text-to-image alignment and better aesthetic metrics, demonstrating higher generation quality and utility.
Unified LoRA. In addition to the different steps of LoRAs proposed above, we note that our one-step LoRA can be considered as a unified approach, since it is able to reason about different number of steps (e.g. 1,2,4,8 as shown in Sec. C.3) and consistently generate high-quality results under the effect of consistency distillation. For completeness, Tab. 4 also presents the quantitative results of different steps when applying the 1-Step unified LoRA. We can observe that there is no difference in image-text matching between different steps as the CLIPScore evaluates, which means that user prompts are well adhered to. And as the other metrics show, the aesthetics rise slightly as the step increases, which is as expected after all the user can choose based on the needs for efficiency. This would be of great convenience and practicality in real-world deployment scenarios, since generally only one model can be loaded per instance.
Compatibility with ControlNet. Sec. C.4 shows that our models are also compatible with ControlNet [40]. We test the one-step unified SD15 and SDXL LoRAs on the scribble [4] and canny [1] control images, respectively. And we can observe the conditions are well followed and the consistency of our unified LoRAs can still be demonstrated, where the quality of generated images under different inference steps are always guaranteed.
Compatibility with Base Model. Sec. C.5 shows that our LoRAs can be applied to different base models. Specifically, we conduct experiments on anime [2], realistic [3] and artstyle [5] base models. The results demonstrate that our method has a wide range of applications, and the lightweight LoRA also significantly reduces the cost of acceleration.
5 Conclusion
We propose Hyper-SD, a unified framework that maximizes the few-step generation capacity of diffusion models, achieving new SOTA performance based on SDXL and SD15. By employing trajectory-segmented consistency distillation, we enhanced the trajectory preservation ability during distillation, approaching the generation proficiency of the original model. Then, human feedback learning and variational score distillation stimulated the potential for few-step inference, resulting in a more optimal and efficient trajectory for generating models. We have open-sourced LoRAs for SDXL and SD15 from 1 to 8 steps inference, along with a dedicated one-step SDXL model, aiming to further propel the development of generative AI community. More discussions refer to Appendix D.
Appendix A Preliminaries of Diffusion Model
$\mathrm{d}x_{t}=\mu(x_{t},t)\mathrm{d}t+\sigma(t)\mathrm{d}w_{t},$ (16)
Appendix B Pseudo Code of TSCD
@@
Appendix C Qualitative Results
C.1 SD15 Architecture with LoRA training
@@ C.2 SDXL Architecture with UNet training
C.3 Unified LoRA
@@
C.4 Compatibility with ControlNet
C.5 Compatibility with Base Model
@@
Appendix D Discussion and Limitation